OpenStudy (anonymous):

Solve for x. -3x - 8 + 6x = 7

OpenStudy (callisto):

-3x - 8 + 6x = 7 -3x + 6x -8 =7 3x -8 =7 3x - 8 +8 = 7+8 3x = 15 x=?

OpenStudy (anonymous):

5